OpenStudy (anonymous):

Use the Distributive Property and simplify the expression. 9 + 4(2x – 1) + 8 2x + 20 6x + 20 8x + 13 8x + 16

OpenStudy (jojo4eva):

8x +13

OpenStudy (jojo4eva):

4*2x=8x and u get -4+8+9=13

OpenStudy (jojo4eva):

got it chica

OpenStudy (anonymous):

Solve. z + 17 = 9 (Points : 4) –9 –8 8 9

OpenStudy (anonymous):

may u explain more about how u got the answer

OpenStudy (jojo4eva):

subtract 17 from both sides and u get -8

OpenStudy (jojo4eva):

For the first one i used the distributive property and multiplied 4x2x=8x and multiplied 4 by -1 and got -4 so the equation is then 9+8x-4+8. Compare like terms and you get 8x+13

OpenStudy (jojo4eva):

Got it?

OpenStudy (anonymous):

yup i got it 100%

OpenStudy (jojo4eva):

coolio

OpenStudy (anonymous):

Solve. 2|s| = 18 s = –9 only s = 9 only s = –18 or s = 18 s = –9 or s = 9

OpenStudy (anonymous):

lol i say coolio to ha

OpenStudy (jojo4eva):

ya habit lol

OpenStudy (anonymous):

ok i believe this one would be B but then i seen D

OpenStudy (anonymous):

lol ya its a habit for me to i say it lot lately

OpenStudy (anonymous):

which one would this one be Solve. 2|s| = 18 s = –9 only s = 9 only s = –18 or s = 18 s = –9 or s = 9

OpenStudy (jojo4eva):

its would be the last 1 because remember absolute value automatically makes the value positive

OpenStudy (anonymous):

oh ok so both a positive and negative

OpenStudy (jojo4eva):

yep because the absolute value takes away the negative making it an absolute value

OpenStudy (jojo4eva):

ANy other questions?

OpenStudy (anonymous):

Solve. 2x + 6 – 6x = 30 –6 –3 3 6

OpenStudy (jojo4eva):

add the xs and u get -4x and then subtract 6 from both sides and you get -4x=24. divide both sides by -4 and x=6

OpenStudy (jojo4eva):

Any other questions lol? (knowing that you probably have more lol)

OpenStudy (anonymous):

lol this is it

OpenStudy (jojo4eva):

i believe its the last one

OpenStudy (jojo4eva):

the last answer

OpenStudy (anonymous):

do u know how to do linear equations

OpenStudy (jojo4eva):

ya but i kinda got to go lol i have choir rehearsal at my church and im playing for them so can i answer them for you later one?

OpenStudy (anonymous):

Subtract: –7 – (–5) –12 –2 2 12 like this

OpenStudy (anonymous):

i got -2

OpenStudy (anonymous):

@jojo4eva

OpenStudy (jojo4eva):

that is correct